### v1

# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2024-272 LHCb-PAPER-2024-041

EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)
CERN-EP-2024-272
LHCb-PAPER-2024-041
November 7, 2024
Measurement of the to cross-section ratio as a function of centrality in PbPb collisions at
LHCb collaboration†††Authors are listed at the end of this paper.
The dissociation of quarkonium states with different binding energies produced in heavy-ion collisions is a powerful probe for investigating the formation and properties of the quark-gluon plasma. The ratio of production cross-sections of and mesons times the ratio of their branching fractions into the dimuon final state is measured as a function of centrality using data collected by the LHCb detector in PbPb collisions at = 5.02 TeV. The measured ratio shows no dependence on the collision centrality, and is compared to the latest theory predictions and to the recent measurements in literature.
Submitted to JHEP
© 2024 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
1 Introduction
The production of charmonia, in particular of the and states, has long been considered a key probe for understanding the properties of the quark-gluon plasma (QGP). The idea that quarkonium suppression is an indicator of QGP formation was introduced by T. Matsui and H. Satz, who proposed that in deconfined nuclear matter, pairs become unbound under the effect of colour screening [1]. Following their theory, quarkonium states, and in particular those that are loosely bound (Fig. 1), dissociate in the hot and dense medium produced in heavy-ion collisions. Experimentally, the relative yields between different quarkonium states are thus expected to change with the centrality of the collision, the latter being closely related to the temperature of the medium.
Early results from the NA50 collaboration at the SPS showed suppression in PbPb collisions [2, 3]. These results were extensively debated but were not a conclusive signature of QGP formation as other competing mechanisms can also contribute to quarkonium suppression, including cold nuclear matter (CNM) effects such as shadowing and nuclear absorption, and final-state interactions with comoving particles. Further studies by the E866 [4], PHENIX [5], ALICE [6], CMS [7] and LHCb [8, 9] experiments showed that the suppression is observed even in smaller collision systems like proton-nucleus (A), known to be less prone to colour-screening effects. In particular, the suppression of production relative to that of states in A collisions suggests that comovers, rather than QGP effects alone, are important in understanding quarkonium behaviour in high-energy collisions. Theoretical models accounting for these effects have been developed to provide a more comprehensive interpretation of the data [10].
In addition, production in heavy-ion collisions was extensively studied by RHIC and LHC experiments which played a crucial role in understanding quarkonium suppression in a hot and dense medium. Results from the PHENIX [11] and STAR [12] experiments demonstrated significant suppression of yields in AuAu collisions, while also providing insight into the competing effects of QGP formation and CNM contributions. These findings were then complemented by the ALICE [13] and CMS [14] experiments which studied charmonium production in a variety of collision systems including PbPb, and also by the LHCb collaboration, which focused on the different production mechanisms [15]. Together, these studies have advanced the understanding of the interplay between cold and hot nuclear matter effects, motivating new investigations of the properties of the medium through the relative suppression of quarkonium states.
In this paper, the relative production of to states, both decaying to two muons, is studied as a function of centrality using the 2018 PbPb collision data collected at a nucleon-nucleon centre-of-mass energy of by the LHCb experiment, corresponding to an integrated luminosity of around [18]. The ratio of cross-sections () of the two charmonium states multiplied by their branching fractions () in the dimuon channel is measured as
where and are the signal yields and efficiencies, respectively.
2 The LHCb detector and data selection
The LHCb detector [19, 20] is a single-arm forward spectrometer covering the pseudorapidity range 2 5, designed for the study of particles containing or quarks. The detector includes a high-precision tracking system consisting of a silicon-strip vertex detector (VELO) surrounding the interaction region [21], a large-area silicon-strip detector located upstream of a dipole magnet with a bending power of about , and three stations of silicon-strip detectors and straw drift tubes [22] placed downstream of the magnet. The tracking system provides a measurement of the momentum, , of charged particles with a relative uncertainty that varies from 0.5% at low momentum to 1.0% at 200. The minimum distance of a track to a primary vertex (PV), the impact parameter, is measured with a resolution of , where is the component of the momentum transverse to the beam, in . Different types of charged hadrons are distinguished using information from two ring-imaging Cherenkov detectors (RICH) [23]. Photons, electrons and hadrons are identified by a calorimeter system consisting of scintillating-pad and preshower detectors, an electromagnetic and a hadronic calorimeter (ECAL and HCAL). Muons are identified by a system composed of alternating layers of iron and multiwire proportional chambers [24].
The online event selection is performed by a trigger [25], which consists of a hardware stage, based on information from the calorimeter and muon systems, followed by a software stage, which applies a full event reconstruction. At the hardware level, events are required to have a muon with high . Events are selected at software level if the number of clusters in the VELO, , satisfies or if and two muons with are identified. Moreover, to suppress contamination from Pb-gas interactions,111Neon gas was injected in the beam pipe near the interaction point, using the LHCb fixed-target SMOG system [18], concurrently with PbPb collisions. a minimum of 15 VELO tracks in the backward direction, , is required, along with the condition that the PV must be close to the beam collision point. The and candidates are reconstructed in the dimuon final state, satisfying selection requirements on transverse momentum, , and rapidity, . The two muons are both required to have greater then and to satisfy particle identification (PID) requirements based on information from the RICH, calorimeter and muon systems.
3 Yields in centrality intervals
The charmonium yields are determined in intervals of the collision centrality. Centrality intervals are defined as percentiles of the total inelastic hadronic PbPb cross-section and are associated with the impact parameter of the colliding particles and the mean number of participating nucleons, [26]. A higher centrality percentile corresponds to more peripheral collisions with a larger impact parameter, while a lower centrality percentile corresponds to more central collisions with a smaller impact parameter. To estimate these geometric quantities, the Glauber Monte Carlo (GMC) model adapted to LHCb conditions is used [26, 27]. The impact parameter cannot be measured directly and is instead estimated using the total energy deposited in the ECAL, with which it scales linearly. The method is based on a binned fit of the total ECAL energy in simulated minimum-bias PbPb interactions where the GMC model is used and the experimental conditions of the signal sample are reproduced. This fit establishes a per-event dependence of and the collision centrality in terms of the energy deposited in the ECAL. The centrality intervals used are defined in Table 1. The data used to determine the ratio are limited to centrality values down to about due to the inability of the detector to cope with the increasing occupancy.
The charmonium yields, and , are extracted by an unbinned maximum-likelihood fit to the dimuon invariant-mass distribution in each centrality interval where the signal is modelled by a Crystal Ball function [28] with tail parameters fixed from simulation. A Gaussian function is used to describe the signal; the difference between the peak positions of the and signals is fixed to the difference of the known masses of the two states [29]. The background is described by an exponential function.
Centrality (%) | 90–100 | 80–90 | 70–80 | 60–70 |
2.4–5.5 | 5.5–13.0 | 13.0–26.5 | 26.5–48.0 | |
Figure 2 shows the dimuon invariant-mass distribution in each centrality interval compared to the fit results. The and yields extracted by the fit are shown in Table 1.
4 Simulation and efficiencies
The total signal selection efficiencies in Eq. 1 are determined using simulation. The simulated decays are reconstructed and analysed using the same software tools as those used to process the data. In the simulation, and particles are generated using Pythia [30] and embedded into minimum-bias PbPb collisions from the EPOS event generator [31]. The decays of unstable particles are described by EvtGen [32], in which final-state radiation is generated using PHOTOS [33]. The interaction of the generated particles with the detector and its response are implemented using the GEANT4 toolkit [34, 35] adapted to LHCb [36]. After reconstruction, the simulated events are weighted to improve agreement with data. The correction weights are obtained as a function of and from the background-subtracted data and simulation sample of candidates. The total efficiencies include the geometric acceptance of the detector, the candidate reconstruction and selection, the particle identification and the effect of the trigger, which are estimated in each centrality interval, using the weighted simulation.
The acceptance efficiency, defined as the fraction of decays with both muons entering the acceptance of the LHCb spectrometer, is determined using a simulation sample produced without any requirement on the kinematics of the decay products. This efficiency is found to be independent of centrality. The combined reconstruction and selection efficiency is calculated as the fraction of generated events in the acceptance that pass the selection requirements other than those related to PID and the trigger. The PID efficiency for the two muons is estimated from dedicated calibration data samples, and computed as a function of the muon kinematics. Specifically, the muon PID efficiency is determined in intervals of muon and , showing a smooth dependence up to the edge of the multiplicity region at . In each muon bin, it is then extrapolated via a quadratic fit function to values up to 10000, to cover the multiplicity region of the measurement reported here using PbPb collisions. Finally, the trigger efficiency is computed as the fraction of simulated signal particles that pass the trigger requirement.
5 Systematic uncertainties
Several sources of systematic uncertainty associated with the determination of the signal and background yields and with the evaluation of the efficiencies are considered.
The systematic uncertainty arising from the signal parametrisation consists of two contributions. The first is associated with the choice of the fixed parameters of the signal fit function. It is estimated by varying the parameters within their uncertainties obtained in simulation. The second contribution is estimated by repeating the signal yield extraction with an alternative parametrisation of the signal using the convolution of a Crystal Ball and a Gaussian function where both widths are free to vary. Compared to the large statistical uncertainty affecting the yield, the systematic uncertainty arising from the fit model is negligibly small and ignored. The uncertainty in background modelling is assessed by comparing the results from the baseline fit to those obtained using a second-order polynomial instead of the exponential function.
The systematic uncertainty arising from the evaluation of the efficiencies consists of different contributions. The first contribution is related to the weighting procedure applied to the simulation, and is taken into consideration by randomly varying each weight within its uncertainty and estimating the total efficiencies again. Using 50 different sets of generated weights, the uncertainty considered is the relative variation of the obtained efficiencies. Another contribution pertains to the estimation of the PID efficiency. An alternative approach is used, where the extrapolation to the high-multiplicity region that is not covered by the calibration sample is performed with a linear function instead of a second-order polynomial. The limited size of the simulation sample used to determine the efficiencies is also a source of systematic uncertainty, and is taken into account using the statistical uncertainties of the efficiencies determined from simulation. Finally, a systematic uncertainty due to the trigger efficiency is assigned by using a data-driven method where the efficiency is determined from events triggered independently of the signal muons [37]. The total systematic uncertainty is computed as the sum in quadrature of each contribution shown in Table 2.
Source | Uncertainty (%) |
Signal fit model | 0.5–2.2 |
Background fit model | 0.5–6.3 |
Weighting procedure | 0.1–0.2 |
PID efficiency extrapolation | 0.1–2.8 |
Simulation sample size | 0.5–1.0 |
Trigger efficiency estimation | 0.1 |
6 Results
The results in different centrality intervals are listed in Table 3 and are presented as a function of the number of participating nucleons in Figs. 3 and 4, where they are compared with some of the recently published measurements and with theory predictions, respectively. In particular, Fig. 3 compares these results to the ALICE measurement [38], performed in PbPb collisions at the same centre-of-mass energy and at similar rapidities and transverse momenta but covering different centrality region. The two measurements are in agreement within uncertainties in the overlapping region, showing no dependency on . The LHCb [39] and Pb [40] measurements are reported in the same figure showing also agreement with the results reported here. The ratio measured in these studies is integrated over the kinematic range, and is associated with a number of participating nucleons of = 2 (8) for the (Pb) measurement, as estimated using the GMC model [27].
Figure 4 shows a comparison of these results with two recent theoretical calculations. In the statistical hadronization model (SHMc) [41, 42], it is assumed that, near the phase boundary between the QGP at high temperatures and the confined hadronic matter at lower temperatures, a fireball is formed during collisions, which is near thermal equilibrium. The TAMU model [43, 44] is a transport model developed in a nonperturbative hydro-Langevin-RRM framework that includes collisional energy loss and heavy-quark diffusion in the medium. When compared to the data reported in this article, the SHMc model, although showing no dependency of the ratio on the centrality, generally underestimates its value. On the other hand, the TAMU model predictions are in better agreement with the experimental result, especially in the lower multiplicity intervals. A slight disagreement is observed in the highest multiplicity interval, where the data differ from the two theoretical predictions and the ALICE measurement by approximately 1.5 standard deviations.
Centrality (%) | ||
90–100 | 2.40–5.50 | |
80–90 | 5.50–13.0 | |
70–80 | 13.0–26.5 | |
60–70 | 26.5–48.0 |
7 Conclusions
The to production cross-section ratio in PbPb collisions at a centre-of-mass energy of corresponding to an integrated luminosity of around is measured as a function of event centrality. No significant dependence on the centrality is found. The results are also compared to the latest theory predictions, with which they agree within 1.5 standard deviations or better, with a slight tension in the lowest centrality interval. The SHMc calculations tend to underestimate the data over the whole range; the TAMU model achieves a better description. However, the current experimental precision precludes any firm conclusion. This study is a first step towards a deeper understanding of quarkonium production in PbPb collisions at midcentrality that can be achieved by additional measurements with the upgraded LHCb detector in Run 3 [45].
Acknowledgements
We thank Anton Andronic, Ralf Rapp and Biao Gang for providing the theoretical predictions for our measurement. We express our gratitude to our colleagues in the CERN accelerator departments for the excellent performance of the LHC. We thank the technical and administrative staff at the LHCb institutes. We acknowledge support from CERN and from the national agencies: CAPES, CNPq, FAPERJ and FINEP (Brazil); MOST and NSFC (China); CNRS/IN2P3 (France); BMBF, DFG and MPG (Germany); INFN (Italy); NWO (Netherlands); MNiSW and NCN (Poland); MCID/IFA (Romania); MICIU and AEI (Spain); SNSF and SER (Switzerland); NASU (Ukraine); STFC (United Kingdom); DOE NP and NSF (USA). We acknowledge the computing resources that are provided by CERN, IN2P3 (France), KIT and DESY (Germany), INFN (Italy), SURF (Netherlands), PIC (Spain), GridPP (United Kingdom), CSCS (Switzerland), IFIN-HH (Romania), CBPF (Brazil), and Polish WLCG (Poland). We are indebted to the communities behind the multiple open-source software packages on which we depend. Individual groups or members have received support from ARC and ARDC (Australia); Key Research Program of Frontier Sciences of CAS, CAS PIFI, CAS CCEPP, Fundamental Research Funds for the Central Universities, and Sci. & Tech. Program of Guangzhou (China); Minciencias (Colombia); EPLANET, Marie Skłodowska-Curie Actions, ERC and NextGenerationEU (European Union); A*MIDEX, ANR, IPhU and Labex P2IO, and Région Auvergne-Rhône-Alpes (France); AvH Foundation (Germany); ICSC (Italy); Severo Ochoa and María de Maeztu Units of Excellence, GVA, XuntaGal, GENCAT, InTalent-Inditex and Prog. Atracción Talento CM (Spain); SRC (Sweden); the Leverhulme Trust, the Royal Society and UKRI (United Kingdom).
F. Abudinén57
T. Ackernley61
A. A. Adefisoye69
B. Adeva47
M. Adinolfi55
P. Adlarson82
C. Agapopoulou14
C.A. Aidala83
Z. Ajaltouni11,
S. Akar66
K. Akiba38
P. Albicocco28
J. Albrecht19
F. Alessio49
M. Alexander60
Z. Aliouche63
P. Alvarez Cartelle56
R. Amalric16
S. Amato3
J.L. Amey55
Y. Amhis14
L. An6
L. Anderlini27
M. Andersson51
A. Andreianov44
P. Andreola51
M. Andreotti26
D. Andreou69
A. Anelli31,n
D. Ao7
F. Archilli37,t
M. Argenton26
S. Arguedas Cuendis9,49
A. Artamonov44
M. Artuso69
E. Aslanides13
R. Ataíde Da Silva50
M. Atzeni65
B. Audurier12
D. Bacher64
I. Bachiller Perea10
S. Bachmann22
M. Bachmayer50
J.J. Back57
P. Baladron Rodriguez47
V. Balagura15
A. Balboni26
W. Baldini26
L. Balzani19
H. Bao7
J. Baptista de Souza Leite61
C. Barbero Pretel47,12
M. Barbetti27
I. R. Barbosa70
R.J. Barlow63
M. Barnyakov25
S. Barsuk14
W. Barter59
M. Bartolini56
J. Bartz69
J.M. Basels17
S. Bashir40
G. Bassi35,q
B. Batsukh5
P. B. Battista14,
A. Bay50
A. Beck57
M. Becker19
F. Bedeschi35
I.B. Bediaga2
N. A. Behling19
S. Belin47
K. Belous44
I. Belov29
I. Belyaev36
G. Benane13
G. Bencivenni28
E. Ben-Haim16
A. Berezhnoy44
R. Bernet51
S. Bernet Andres45
A. Bertolin33
C. Betancourt51
F. Betti59
J. Bex56
Ia. Bezshyiko51
J. Bhom41
M.S. Bieker19
N.V. Biesuz26
P. Billoir16
A. Biolchini38
M. Birch62
F.C.R. Bishop10
A. Bitadze63
A. Bizzeti
T. Blake57
F. Blanc50
J.E. Blank19
S. Blusk69
V. Bocharnikov44
J.A. Boelhauve19
O. Boente Garcia15
T. Boettcher66
A. Bohare59
A. Boldyrev44
C.S. Bolognani79
R. Bolzonella26,k
R. B. Bonacci1
N. Bondar44
A. Bordelius49
F. Borgato33,o
S. Borghi63
M. Borsato31,n
J.T. Borsuk41
S.A. Bouchiba50
M. Bovill64
T.J.V. Bowcock61
A. Boyer49
C. Bozzi26
A. Brea Rodriguez50
N. Breer19
J. Brodzicka41
A. Brossa Gonzalo47,†
J. Brown61
D. Brundu32
E. Buchanan59,
A. Buonaura51
L. Buonincontri33,o
A.T. Burke63
C. Burr49
J.S. Butter56
J. Buytaert49
W. Byczynski49
S. Cadeddu32
H. Cai74,
A. C. Caillet16,
R. Calabrese26,k
S. Calderon Ramirez9
L. Calefice46
S. Cali28
M. Calvi31,n
M. Calvo Gomez45
P. Camargo Magalhaes2,x
J. I. Cambon Bouzas47
P. Campana28
D.H. Campora Perez79
A.F. Campoverde Quezada7
S. Capelli31
L. Capriotti26
R. Caravaca-Mora9
A. Carbone25,i
L. Carcedo Salgado47
R. Cardinale29,l
A. Cardini32
P. Carniti31,n
L. Carus22,
A. Casais Vidal65
R. Caspary22
G. Casse61
M. Cattaneo49
G. Cavallero26,49
V. Cavallini26,k
S. Celani22
D. Cervenkov64
S. Cesare30,m
A.J. Chadwick61
I. Chahrour83
M. Charles16
Ph. Charpentier49
E. Chatzianagnostou38
M. Chefdeville10
C. Chen13
S. Chen5
Z. Chen7
A. Chernov41
S. Chernyshenko53
X. Chiotopoulos79
V. Chobanova81
S. Cholak50
M. Chrzaszcz41
A. Chubykin44
V. Chulikov28
P. Ciambrone28
X. Cid Vidal47
G. Ciezarek49
P. Cifra49
P.E.L. Clarke59
M. Clemencic49
H.V. Cliff56
J. Closier49
C. Cocha Toapaxi22
V. Coco49
J. Cogan13
E. Cogneras11
L. Cojocariu43
S. Collaviti50
P. Collins49
T. Colombo49
M. C. Colonna19
A. Comerma-Montells46
L. Congedo24
A. Contu32
N. Cooke60
I. Corredoira 47
A. Correia16
G. Corti49
J.J. Cottee Meldrum55,
B. Couturier49
D.C. Craik51
M. Cruz Torres2,f
E. Curras Rivera50
R. Currie59
C.L. Da Silva68
S. Dadabaev44
L. Dai71
X. Dai6
E. Dall’Occo49
J. Dalseno47
C. D’Ambrosio49
J. Daniel11
A. Danilina44
P. d’Argent24
G. Darze3,
A. Davidson57
J.E. Davies63
A. Davis63
O. De Aguiar Francisco63
C. De Angelis32,j
F. De Benedetti49
J. de Boer38
K. De Bruyn78
S. De Capua63
M. De Cian22
U. De Freitas Carneiro Da Graca2,a
E. De Lucia28
J.M. De Miranda2
L. De Paula3
M. De Serio24,g
P. De Simone28
F. De Vellis19
J.A. de Vries79
F. Debernardis24
D. Decamp10
V. Dedu13
S. Dekkers1
L. Del Buono16
B. Delaney65
H.-P. Dembinski19
J. Deng8
V. Denysenko51
O. Deschamps11
F. Dettori32,j
B. Dey77
P. Di Nezza28
I. Diachkov44
S. Didenko44
S. Ding69
L. Dittmann22
V. Dobishuk53
A. D. Docheva60
C. Dong4,b
A.M. Donohoe23
F. Dordei32
A.C. dos Reis2
A. D. Dowling69
W. Duan72
P. Duda80
M.W. Dudek41
L. Dufour49
V. Duk34
P. Durante49
M. M. Duras80
J.M. Durham68
O. D. Durmus77
A. Dziurda41
A. Dzyuba44
S. Easo58
E. Eckstein18,
U. Egede1
A. Egorychev44
V. Egorychev44
S. Eisenhardt59
E. Ejopu63
L. Eklund82
M. Elashri66
J. Ellbracht19
S. Ely62
A. Ene43
J. Eschle69
S. Esen22
T. Evans63
F. Fabiano32,j
L.N. Falcao2
Y. Fan7
B. Fang7
L. Fantini34,p,49
M. Faria50
K. Farmer59
D. Fazzini31,n
L. Felkowski80
M. Feng5,7
M. Feo19
A. Fernandez Casani48
M. Fernandez Gomez47
A.D. Fernez67
F. Ferrari25
F. Ferreira Rodrigues3
M. Ferrillo51
M. Ferro-Luzzi49
S. Filippov44
R.A. Fini24
M. Fiorini26,k
M. Firlej40
K.L. Fischer64
D.S. Fitzgerald83
C. Fitzpatrick63
T. Fiutowski40
F. Fleuret15
M. Fontana25
L. F. Foreman63
R. Forty49
D. Foulds-Holt56
V. Franco Lima3
M. Franco Sevilla67
M. Frank49
E. Franzoso26,k
G. Frau63
C. Frei49
D.A. Friday63
J. Fu7
Q. Fuehring19,56
Y. Fujii1
T. Fulghesu16
E. Gabriel38
G. Galati24
M.D. Galati38
A. Gallas Torreira47
D. Galli25,i
S. Gambetta59
M. Gandelman3
P. Gandini30
B. Ganie63
H. Gao7
R. Gao64
T.Q. Gao56
Y. Gao8
Y. Gao6
Y. Gao8,
L.M. Garcia Martin50
P. Garcia Moreno46
J. García Pardiñas49
P. Gardner67
K. G. Garg8
L. Garrido46
C. Gaspar49
R.E. Geertsema38
L.L. Gerken19
E. Gersabeck63
M. Gersabeck20
T. Gershon57
S. G. Ghizzo29,
Z. Ghorbanimoghaddam55,
L. Giambastiani33,o
F. I. Giasemis16,e
V. Gibson56
H.K. Giemza42
A.L. Gilman64
M. Giovannetti28
A. Gioventù46
L. Girardey63
P. Gironella Gironell46
C. Giugliano26,k
M.A. Giza41
E.L. Gkougkousis62
F.C. Glaser14,22
V.V. Gligorov16,49
C. Göbel70
E. Golobardes45
D. Golubkov44
A. Golutvin62,44,49
S. Gomez Fernandez46
W. Gomulka40,
F. Goncalves Abrantes64
M. Goncerz41
G. Gong4,b
J. A. Gooding19
I.V. Gorelov44
C. Gotti31
J.P. Grabowski18
L.A. Granado Cardoso49
E. Graugés46
E. Graverini50,r
L. Grazette57
G. Graziani
A. T. Grecu43
L.M. Greeven38
N.A. Grieser66
L. Grillo60
S. Gromov44
C. Gu15
M. Guarise26
L. Guerry11
M. Guittiere14
V. Guliaeva44
P. A. Günther22
A.-K. Guseinov50
E. Gushchin44
Y. Guz6,44,49
T. Gys49
K. Habermann18
T. Hadavizadeh1
C. Hadjivasiliou67
G. Haefeli50
C. Haen49
M. Hajheidari49,
G. Hallett57
M.M. Halvorsen49
P.M. Hamilton67
J. Hammerich61
Q. Han8
X. Han22,49
S. Hansmann-Menzemer22
L. Hao7
N. Harnew64
T. H. Harris1
M. Hartmann14
S. Hashmi40
J. He7,c
F. Hemmer49
C. Henderson66
R.D.L. Henderson1,57
A.M. Hennequin49
K. Hennessy61
L. Henry50
J. Herd62
P. Herrero Gascon22
J. Heuel17
A. Hicheur3
G. Hijano Mendizabal51,
J. Horswill63
R. Hou8
Y. Hou11
N. Howarth61,
J. Hu72
W. Hu6
X. Hu4,b
W. Huang7
W. Hulsbergen38
R.J. Hunter57
M. Hushchyn44
D. Hutchcroft61
M. Idzik40
D. Ilin44
P. Ilten66
A. Inglessi44
A. Iniukhin44
A. Ishteev44
K. Ivshin44
R. Jacobsson49
H. Jage17
S.J. Jaimes Elles75,49,48
S. Jakobsen49
E. Jans38
B.K. Jashal48
A. Jawahery67,49
V. Jevtic19
E. Jiang67
X. Jiang5,7
Y. Jiang7
Y. J. Jiang6
M. John64
A. John Rubesh Rajan23
D. Johnson54
C.R. Jones56
T.P. Jones57
S. Joshi42
B. Jost49
J. Juan Castella56
N. Jurik49
I. Juszczak41
D. Kaminaris50
S. Kandybei52
M. Kane59
Y. Kang4,b
C. Kar11
M. Karacson49
D. Karpenkov44
A. Kauniskangas50
J.W. Kautz66
M.K. Kazanecki41,
F. Keizer49
M. Kenzie56
T. Ketel38
B. Khanji69
A. Kharisova44
S. Kholodenko35,49
G. Khreich14
T. Kirn17
V.S. Kirsebom31,n
O. Kitouni65
S. Klaver39
N. Kleijne35,q
K. Klimaszewski42
M.R. Kmiec42
S. Koliiev53
L. Kolk19
A. Konoplyannikov44
P. Kopciewicz40,49
P. Koppenburg38
M. Korolev44
I. Kostiuk38
O. Kot53,
S. Kotriakhova
A. Kozachuk44
P. Kravchenko44
L. Kravchuk44
M. Kreps57
P. Krokovny44
W. Krupa69
W. Krzemien42
O.K. Kshyvanskyi53,
S. Kubis80
M. Kucharczyk41
V. Kudryavtsev44
E. Kulikova44
A. Kupsc82
B. K. Kutsenko13
D. Lacarrere49
P. Laguarta Gonzalez46
A. Lai32
A. Lampis32
D. Lancierini56
C. Landesa Gomez47
J.J. Lane1
R. Lane55
G. Lanfranchi28
C. Langenbruch22
J. Langer19
O. Lantwin44
T. Latham57
F. Lazzari35,r
C. Lazzeroni54
R. Le Gac13
H. Lee61
R. Lefèvre11
A. Leflat44
S. Legotin44
M. Lehuraux57
E. Lemos Cid49
O. Leroy13
T. Lesiak41
E. Lesser49,
B. Leverington22
A. Li4,b
C. Li13
H. Li72
K. Li8
L. Li63
M. Li8,
P. Li7
P.-R. Li73
Q. Li5,7
S. Li8
T. Li5,d
T. Li72
Y. Li8,
Y. Li5
Z. Lian4,b
X. Liang69
S. Libralon48
C. Lin7
T. Lin58
R. Lindner49
H. Linton62
V. Lisovskyi50
R. Litvinov32,49
F. L. Liu1
G. Liu72
K. Liu73
S. Liu5,7
W. Liu8,
Y. Liu59
Y. Liu73,
Y. L. Liu62
A. Lobo Salvia46
A. Loi32
T. Long56
J.H. Lopes3
A. Lopez Huertas46
S. López Soliño47
Q. Lu15
C. Lucarelli27
D. Lucchesi33,o
M. Lucio Martinez79
V. Lukashenko38,53
Y. Luo6
A. Lupato33,h
E. Luppi26,k
K. Lynch23
X.-R. Lyu7
G. M. Ma4,b
S. Maccolini19
F. Machefert14
F. Maciuc43
B. Mack69
I. Mackay64
L. M. Mackey69
L.R. Madhan Mohan56
M. J. Madurai54
A. Maevskiy44
D. Magdalinski38
D. Maisuzenko44
M.W. Majewski40,
J.J. Malczewski41
S. Malde64
L. Malentacca49,
A. Malinin44
T. Maltsev44
G. Manca32,j
G. Mancinelli13
C. Mancuso30,14,m
R. Manera Escalero46
F. M. Manganella37,
D. Manuzzi25
D. Marangotto30,m
J.F. Marchand10
R. Marchevski50
U. Marconi25
E. Mariani16,
S. Mariani49
C. Marin Benito46,49
J. Marks22
A.M. Marshall55
L. Martel64
G. Martelli34,p
G. Martellotti36
L. Martinazzoli49
M. Martinelli31,n
D. Martinez Gomez78
D. Martinez Santos81
F. Martinez Vidal48
A. Martorell i Granollers45
A. Massafferri2
R. Matev49
A. Mathad49
V. Matiunin44
C. Matteuzzi69
K.R. Mattioli15
A. Mauri62
E. Maurice15
J. Mauricio46
P. Mayencourt50
J. Mazorra de Cos48
M. Mazurek42
M. McCann62
L. Mcconnell23
T.H. McGrath63
N.T. McHugh60
A. McNab63
R. McNulty23
B. Meadows66
G. Meier19
D. Melnychuk42
F. M. Meng4,b
M. Merk38,79
A. Merli50
L. Meyer Garcia67
D. Miao5,7
H. Miao7
M. Mikhasenko76
D.A. Milanes75
A. Minotti31,n
E. Minucci28
T. Miralles11
B. Mitreska19
D.S. Mitzel19
A. Modak58
R.A. Mohammed64
R.D. Moise17
S. Mokhnenko44
E. F. Molina Cardenas83
T. Mombächer49
M. Monk57,1
S. Monteil11
A. Morcillo Gomez47
G. Morello28
M.J. Morello35,q
M.P. Morgenthaler22
J. Moron40
W. Morren38
A.B. Morris49
A.G. Morris13
R. Mountain69
H. Mu4,b
Z. M. Mu6
E. Muhammad57
F. Muheim59
M. Mulder78
K. Müller51
F. Muñoz-Rojas9
R. Murta62
P. Naik61
T. Nakada50
R. Nandakumar58
T. Nanut49
I. Nasteva3
M. Needham59
N. Neri30,m
S. Neubert18
N. Neufeld49
P. Neustroev44,
J. Nicolini19,14
D. Nicotra79
E.M. Niel49
N. Nikitin44
P. Nogarolli3
P. Nogga18,
C. Normand55
J. Novoa Fernandez47
G. Nowak66
C. Nunez83
H. N. Nur60
A. Oblakowska-Mucha40
V. Obraztsov44
T. Oeser17
S. Okamura26,k
A. Okhotnikov44,
O. Okhrimenko53
R. Oldeman32,j
F. Oliva59
M. Olocco19
C.J.G. Onderwater79
R.H. O’Neil59
D. Osthues19,
J.M. Otalora Goicochea3
P. Owen51
A. Oyanguren48
O. Ozcelik59
F. Paciolla35,u
A. Padee42
K.O. Padeken18
B. Pagare57
P.R. Pais22
T. Pajero49
A. Palano24
M. Palutan28
X. Pan4,b
G. Panshin44
L. Paolucci57
A. Papanestis58,49
M. Pappagallo24,g
L.L. Pappalardo26,k
C. Pappenheimer66
C. Parkes63
D. Parmar76
B. Passalacqua26,k
G. Passaleva27
D. Passaro35,q
A. Pastore24
M. Patel62
J. Patoc64
C. Patrignani25,i
A. Paul69
C.J. Pawley79
A. Pellegrino38
J. Peng5,7
M. Pepe Altarelli28
S. Perazzini25
D. Pereima44
H. Pereira Da Costa68
A. Pereiro Castro47
P. Perret11
A. Perrevoort78
A. Perro49
M.J. Peters66,
K. Petridis55
A. Petrolini29,l
J. P. Pfaller66
H. Pham69
L. Pica35,q
M. Piccini34
L. Piccolo32
B. Pietrzyk10
G. Pietrzyk14
D. Pinci36
F. Pisani49
M. Pizzichemi31,n
V. Placinta43
M. Plo Casasus47
T. Poeschl49
F. Polci16,49
M. Poli Lener28
A. Poluektov13
N. Polukhina44
I. Polyakov44
E. Polycarpo3
S. Ponce49
D. Popov7
S. Poslavskii44
K. Prasanth59
C. Prouve81
D. Provenzano32,j
V. Pugatch53
G. Punzi35,r
S. Qasim51
Q. Q. Qian6
W. Qian7
N. Qin4,b
S. Qu4,b
R. Quagliani49
R.I. Rabadan Trejo57
J.H. Rademacker55
M. Rama35
M. Ramírez García83
V. Ramos De Oliveira70
M. Ramos Pernas57
M.S. Rangel3
F. Ratnikov44
G. Raven39
M. Rebollo De Miguel48
F. Redi30,h
J. Reich55
F. Reiss63
Z. Ren7
P.K. Resmi64
R. Ribatti50
G. R. Ricart15,12
D. Riccardi35,q
S. Ricciardi58
K. Richardson65
M. Richardson-Slipper59
K. Rinnert61
P. Robbe14,49
G. Robertson60
E. Rodrigues61
A. Rodriguez Alvarez46,
E. Rodriguez Fernandez47
J.A. Rodriguez Lopez75
E. Rodriguez Rodriguez47
J. Roensch19,
A. Rogachev44
A. Rogovskiy58
D.L. Rolf49
P. Roloff49
V. Romanovskiy66
A. Romero Vidal47
G. Romolini26
F. Ronchetti50
T. Rong6
M. Rotondo28
S. R. Roy22
M.S. Rudolph69
M. Ruiz Diaz22
R.A. Ruiz Fernandez47
J. Ruiz Vidal82,y
A. Ryzhikov44
J. Ryzka40
J. J. Saavedra-Arias9
J.J. Saborido Silva47
R. Sadek15
N. Sagidova44
D. Sahoo77
N. Sahoo54
B. Saitta32,j
M. Salomoni31,n,49
I. Sanderswood48
R. Santacesaria36
C. Santamarina Rios47
M. Santimaria28,49
L. Santoro 2
E. Santovetti37
A. Saputi26,49
D. Saranin44
A. Sarnatskiy78
G. Sarpis59
M. Sarpis63
C. Satriano36,s
A. Satta37
M. Saur6
D. Savrina44
H. Sazak17
F. Sborzacchi49,28
L.G. Scantlebury Smead64
A. Scarabotto19
S. Schael17
S. Scherl61
M. Schiller60
H. Schindler49
M. Schmelling21
B. Schmidt49
S. Schmitt17
H. Schmitz18,
O. Schneider50
A. Schopper49
N. Schulte19
S. Schulte50
M.H. Schune14
R. Schwemmer49
G. Schwering17
B. Sciascia28
A. Sciuccati49
I. Segal76
S. Sellam47
A. Semennikov44
T. Senger51
M. Senghi Soares39
A. Sergi29
N. Serra51
L. Sestini33
A. Seuthe19
Y. Shang6
D.M. Shangase83
M. Shapkin44
R. S. Sharma69
I. Shchemerov44
L. Shchutska50
T. Shears61
L. Shekhtman44
Z. Shen6
S. Sheng5,7
V. Shevchenko44
B. Shi7
Q. Shi7
Y. Shimizu14
E. Shmanin25
R. Shorkin44
J.D. Shupperd69
R. Silva Coutinho69
G. Simi33,o
S. Simone24,g
N. Skidmore57
T. Skwarnicki69
M.W. Slater54
J.C. Smallwood64
E. Smith65
K. Smith68
M. Smith62
A. Snoch38
L. Soares Lavra59
M.D. Sokoloff66
F.J.P. Soler60
A. Solomin44,55
A. Solovev44
I. Solovyev44
N. S. Sommerfeld18,
R. Song1
Y. Song50
Y. Song4,b
Y. S. Song6
F.L. Souza De Almeida69
B. Souza De Paula3
E. Spadaro Norella29
E. Spedicato25
J.G. Speer19
E. Spiridenkov44,
P. Spradlin60
V. Sriskaran49
F. Stagni49
M. Stahl49
S. Stahl49
S. Stanislaus64
E.N. Stein49
O. Steinkamp51
O. Stenyakin44,
H. Stevens19
D. Strekalina44
Y. Su7
F. Suljik64
J. Sun32
L. Sun74
D. Sundfeld2
W. Sutcliffe51,
P.N. Swallow54
K. Swientek40
F. Swystun56
A. Szabelski42
T. Szumlak40
Y. Tan4,b
Y. Tang74
M.D. Tat64
A. Terentev44
F. Terzuoli35,u,49
F. Teubert49
E. Thomas49
D.J.D. Thompson54
H. Tilquin62
V. Tisserand11
S. T’Jampens10
M. Tobin5,49
L. Tomassetti26,k
G. Tonani30,m,49
X. Tong6
D. Torres Machado2
L. Toscano19
D.Y. Tou4,b
C. Trippl45
G. Tuci22
N. Tuning38
L.H. Uecker22
A. Ukleja40
D.J. Unverzagt22
B. Urbach59
E. Ursov44
A. Usachov39
A. Ustyuzhanin44
U. Uwer22
V. Vagnoni25
V. Valcarce Cadenas47
G. Valenti25
N. Valls Canudas49
H. Van Hecke68
E. van Herwijnen62
C.B. Van Hulse47,w
R. Van Laak50
M. van Veghel38
G. Vasquez51
R. Vazquez Gomez46
P. Vazquez Regueiro47
C. Vázquez Sierra47
S. Vecchi26
J.J. Velthuis55
M. Veltri27,v
A. Venkateswaran50
M. Verdoglia32
M. Vesterinen57
D. Vico Benet64
P. V. Vidrier Villalba46,
M. Vieites Diaz49
X. Vilasis-Cardona45
E. Vilella Figueras61
A. Villa25
P. Vincent16
F.C. Volle54
D. vom Bruch13
N. Voropaev44
K. Vos79
C. Vrahas59
J. Wagner19
J. Walsh35
E.J. Walton1,57
G. Wan6
C. Wang22
G. Wang8
J. Wang6
J. Wang5
J. Wang4,b
J. Wang74
M. Wang30
N. W. Wang7
R. Wang55
X. Wang8,
X. Wang72
X. W. Wang62
Y. Wang6
Z. Wang14
Z. Wang4,b
Z. Wang30
J.A. Ward57,1
M. Waterlaat49,
N.K. Watson54
D. Websdale62
Y. Wei6
J. Wendel81
B.D.C. Westhenry55
C. White56
M. Whitehead60
E. Whiter54
A.R. Wiederhold63
D. Wiedner19
G. Wilkinson64
M.K. Wilkinson66
M. Williams65
M. J. Williams49,
M.R.J. Williams59
R. Williams56
Z. Williams55
F.F. Wilson58
M. Winn12,
W. Wislicki42
M. Witek41
L. Witola22
G. Wormser14
S.A. Wotton56
H. Wu69
J. Wu8
X. Wu74
Y. Wu6
Z. Wu7
K. Wyllie49
S. Xian72,
Z. Xiang5
Y. Xie8
A. Xu35
J. Xu7
L. Xu4,b
L. Xu4,b
M. Xu57
Z. Xu49
Z. Xu7
Z. Xu5
K. Yang62
S. Yang7
X. Yang6
Y. Yang29,l
Z. Yang6
V. Yeroshenko14
H. Yeung63
H. Yin8
C. Y. Yu6
J. Yu71
X. Yuan5
Y Yuan5,7
E. Zaffaroni50
M. Zavertyaev21
M. Zdybal41
F. Zenesini25,i
C. Zeng5,7
M. Zeng4,b
C. Zhang6
D. Zhang8
J. Zhang7
L. Zhang4,b
S. Zhang71
S. Zhang64
Y. Zhang6
Y. Z. Zhang4,b
Z. Zhang4,b
Y. Zhao22
A. Zharkova44
A. Zhelezov22
S. Z. Zheng6
X. Z. Zheng4,b
Y. Zheng7
T. Zhou6
X. Zhou8
Y. Zhou7
V. Zhovkovska57
L. Z. Zhu7
X. Zhu4,b
X. Zhu8
V. Zhukov17
J. Zhuo48
Q. Zou5,7
D. Zuliani33,o
G. Zunica50
1School of Physics and Astronomy, Monash University, Melbourne, Australia
2Centro Brasileiro de Pesquisas Físicas (CBPF), Rio de Janeiro, Brazil
3Universidade Federal do Rio de Janeiro (UFRJ), Rio de Janeiro, Brazil
4Department of Engineering Physics, Tsinghua University, Beijing, China, Beijing, China
5Institute Of High Energy Physics (IHEP), Beijing, China
6School of Physics State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing, China
7University of Chinese Academy of Sciences, Beijing, China
8Institute of Particle Physics, Central China Normal University, Wuhan, Hubei, China
9Consejo Nacional de Rectores (CONARE), San Jose, Costa Rica
10Université Savoie Mont Blanc, CNRS, IN2P3-LAPP, Annecy, France
11Université Clermont Auvergne, CNRS/IN2P3, LPC, Clermont-Ferrand, France
12Département de Physique Nucléaire (DPhN), Gif-Sur-Yvette, France
13Aix Marseille Univ, CNRS/IN2P3, CPPM, Marseille, France
14Université Paris-Saclay, CNRS/IN2P3, IJCLab, Orsay, France
15Laboratoire Leprince-Ringuet, CNRS/IN2P3, Ecole Polytechnique, Institut Polytechnique de Paris, Palaiseau, France
16LPNHE, Sorbonne Université, Paris Diderot Sorbonne Paris Cité, CNRS/IN2P3, Paris, France
17I. Physikalisches Institut, RWTH Aachen University, Aachen, Germany
18Universität Bonn - Helmholtz-Institut für Strahlen und Kernphysik, Bonn, Germany
19Fakultät Physik, Technische Universität Dortmund, Dortmund, Germany
20Physikalisches Institut, Albert-Ludwigs-Universität Freiburg, Freiburg, Germany
21Max-Planck-Institut für Kernphysik (MPIK), Heidelberg, Germany
22Physikalisches Institut, Ruprecht-Karls-Universität Heidelberg, Heidelberg, Germany
23School of Physics, University College Dublin, Dublin, Ireland
24INFN Sezione di Bari, Bari, Italy
25INFN Sezione di Bologna, Bologna, Italy
26INFN Sezione di Ferrara, Ferrara, Italy
27INFN Sezione di Firenze, Firenze, Italy
28INFN Laboratori Nazionali di Frascati, Frascati, Italy
29INFN Sezione di Genova, Genova, Italy
30INFN Sezione di Milano, Milano, Italy
31INFN Sezione di Milano-Bicocca, Milano, Italy
32INFN Sezione di Cagliari, Monserrato, Italy
33INFN Sezione di Padova, Padova, Italy
34INFN Sezione di Perugia, Perugia, Italy
35INFN Sezione di Pisa, Pisa, Italy
36INFN Sezione di Roma La Sapienza, Roma, Italy
37INFN Sezione di Roma Tor Vergata, Roma, Italy
38Nikhef National Institute for Subatomic Physics, Amsterdam, Netherlands
39Nikhef National Institute for Subatomic Physics and VU University Amsterdam, Amsterdam, Netherlands
40AGH - University of Krakow, Faculty of Physics and Applied Computer Science, Kraków, Poland
41Henryk Niewodniczanski Institute of Nuclear Physics Polish Academy of Sciences, Kraków, Poland
42National Center for Nuclear Research (NCBJ), Warsaw, Poland
43Horia Hulubei National Institute of Physics and Nuclear Engineering, Bucharest-Magurele, Romania
44Affiliated with an institute covered by a cooperation agreement with CERN
45DS4DS, La Salle, Universitat Ramon Llull, Barcelona, Spain
46ICCUB, Universitat de Barcelona, Barcelona, Spain
47Instituto Galego de Física de Altas Enerxías (IGFAE), Universidade de Santiago de Compostela, Santiago de Compostela, Spain
48Instituto de Fisica Corpuscular, Centro Mixto Universidad de Valencia - CSIC, Valencia, Spain
49European Organization for Nuclear Research (CERN), Geneva, Switzerland
50Institute of Physics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Lausanne, Switzerland
51Physik-Institut, Universität Zürich, Zürich, Switzerland
52NSC Kharkiv Institute of Physics and Technology (NSC KIPT), Kharkiv, Ukraine
53Institute for Nuclear Research of the National Academy of Sciences (KINR), Kyiv, Ukraine
54School of Physics and Astronomy, University of Birmingham, Birmingham, United Kingdom
55H.H. Wills Physics Laboratory, University of Bristol, Bristol, United Kingdom
56Cavendish Laboratory, University of Cambridge, Cambridge, United Kingdom
57Department of Physics, University of Warwick, Coventry, United Kingdom
58STFC Rutherford Appleton Laboratory, Didcot, United Kingdom
59School of Physics and Astronomy, University of Edinburgh, Edinburgh, United Kingdom
60School of Physics and Astronomy, University of Glasgow, Glasgow, United Kingdom
61Oliver Lodge Laboratory, University of Liverpool, Liverpool, United Kingdom
62Imperial College London, London, United Kingdom
63Department of Physics and Astronomy, University of Manchester, Manchester, United Kingdom
64Department of Physics, University of Oxford, Oxford, United Kingdom
65Massachusetts Institute of Technology, Cambridge, MA, United States
66University of Cincinnati, Cincinnati, OH, United States
67University of Maryland, College Park, MD, United States
68Los Alamos National Laboratory (LANL), Los Alamos, NM, United States
69Syracuse University, Syracuse, NY, United States
70Pontifícia Universidade Católica do Rio de Janeiro (PUC-Rio), Rio de Janeiro, Brazil, associated to 3
71School of Physics and Electronics, Hunan University, Changsha City, China, associated to 8
72Guangdong Provincial Key Laboratory of Nuclear Science, Guangdong-Hong Kong Joint Laboratory of Quantum Matter, Institute of Quantum Matter, South China Normal University, Guangzhou, China, associated to 4
73Lanzhou University, Lanzhou, China, associated to 5
74School of Physics and Technology, Wuhan University, Wuhan, China, associated to 4
75Departamento de Fisica , Universidad Nacional de Colombia, Bogota, Colombia, associated to 16
76Ruhr Universitaet Bochum, Fakultaet f. Physik und Astronomie, Bochum, Germany, associated to 19
77Eotvos Lorand University, Budapest, Hungary, associated to 49
78Van Swinderen Institute, University of Groningen, Groningen, Netherlands, associated to 38
79Universiteit Maastricht, Maastricht, Netherlands, associated to 38
80Tadeusz Kosciuszko Cracow University of Technology, Cracow, Poland, associated to 41
81Universidade da Coruña, A Coruna, Spain, associated to 45
82Department of Physics and Astronomy, Uppsala University, Uppsala, Sweden, associated to 60
83University of Michigan, Ann Arbor, MI, United States, associated to 69
aCentro Federal de Educacão Tecnológica Celso Suckow da Fonseca, Rio De Janeiro, Brazil
bCenter for High Energy Physics, Tsinghua University, Beijing, China
cHangzhou Institute for Advanced Study, UCAS, Hangzhou, China
dSchool of Physics and Electronics, Henan University , Kaifeng, China
eLIP6, Sorbonne Université, Paris, France
fUniversidad Nacional Autónoma de Honduras, Tegucigalpa, Honduras
gUniversità di Bari, Bari, Italy
hUniversità di Bergamo, Bergamo, Italy
iUniversità di Bologna, Bologna, Italy
jUniversità di Cagliari, Cagliari, Italy
kUniversità di Ferrara, Ferrara, Italy
lUniversità di Genova, Genova, Italy
mUniversità degli Studi di Milano, Milano, Italy
nUniversità degli Studi di Milano-Bicocca, Milano, Italy
oUniversità di Padova, Padova, Italy
pUniversità di Perugia, Perugia, Italy
qScuola Normale Superiore, Pisa, Italy
rUniversità di Pisa, Pisa, Italy
sUniversità della Basilicata, Potenza, Italy
tUniversità di Roma Tor Vergata, Roma, Italy
uUniversità di Siena, Siena, Italy
vUniversità di Urbino, Urbino, Italy
wUniversidad de Alcalá, Alcalá de Henares , Spain
xFacultad de Ciencias Fisicas, Madrid, Spain
yDepartment of Physics/Division of Particle Physics, Lund, Sweden
†Deceased